# Towards Constraint-aware Learning for Resource Allocation in NFV-enabled Networks

### v1

@@ G.2 Scalability Analysis
Large-scale Network Validation. Similar to previous work (Wang et al., 2023), we generate a random Waxman topology (Waxman, 1988) with 500 nodes and nearly 13,000 links, named WX500. This mimics the modern large-scale cloud cluster, which is more challenging due to the increased complexity of the topology. Additionally, we increase the VNE size distribution to a uniform range from 2 to 20 nodes, and set the arrival rate of VN requests $\eta$ to 0.5. All other simulation and hyperparameter settings remain consistent with those outlined in Section 4.1. To adapt the pretrained CONAL model for WX100 to this larger scenario, we fine-tune it on WX500 via transfer learning. The pretraining on WX100 takes approximately 7.274 hours, and the fine-tuning stage on WX500 consumes an additional 4.326 hours. By leveraging transfer learning, we accelerate the training efficiency of CONAL, facilitating the rapid acquisition of a CONAL model suitable for large-scale scenarios. The results are illustrated in Figure 7. We observe that CONAL consistently outperforms the baselines in terms of both VN_ACR and LT_REV. CONAL also demonstrates superior performance in the LT_R2C, which is only marginally lower than NEA-VNE. Regarding the AVG_ST, NEA-VNE, MCTS-VNE, and PSO-VNE exhibit significantly higher time consumption, whereas CONAL maintains a competitive running efficiency similar to PG-CNN, A3C-GCN, and DDPG-ATT. Despite the increased complexity of the WX500 topology, CONAL maintains a balance between performance and computational efficiency. This analysis demonstrates the scalability and efficiency of CONAL in large-scale network scenarios.
@@
Solving Time Scale Analysis. We investigate solving time across different physical network sizes to evaluate CONAL’s time scalability. Specifically, we increase the size of physical network from 200 to 1,000 nodes with a step of 200 to simulate networks of varying scales. Due to more enough link connectivity of large-scale network system, we set the we set the maximum path length $k$ to 4. Figure 8 shows that as the size increases, NRM-VNE, GRC-VNE and GAL-VNE show the most sightly increased average solving time. In contrast, PSO-VNE and GA-VNE experience a rapid growth in solving time due to their reliance on extensive search. CONAL exhibits the similar trend with A3C-GCN and DDPG-ATT, since they are all based on constructive solution paradigm. As the size of physical network increases, the solution time of these algorithms does not explode, still competitive with MCTS and NEA-VNE. Overall, the solving time of CONAL remains efficient even at larger network scales while offering great performance, making it a viable solution for real-time decision-making in large-scale environments.
G.3 Real-world System Validation
@@
G.4 Hyperparameter Impact Study
@@
The impact of update interval of surrogate policy. This parameter $\mu$ controls the update interval of surrogate policy $\pi^{\prime}$ during training. We vary $\mu$ within the range [1, 5, 10, 20, 30] to explore its impact on key performance metrics, as shown in Figure 9. We observe that increasing $\mu$ initially leads to an improvement in all metrics. This suggests that extremely frequent updates of the surrogate policy make the budget values change rapidly, potentially leading to instability and divergence in the learning process. As $\mu$ increases beyond nearly 10, the improvements across these metrics tend to be stable and show minimal variation. This implies that while a moderate update interval enhances the model’s performance, too-slow updates do not offer a significant further improvement on performance metrics and may even increase computational overhead for training.
The impact of the augment ratio. This parameter determines the proportion of links to be added based on the number of nodes in the network. We change the augment ratio $\epsilon$ within the range [0, 0.5, 1.0, 1.5, 2.0] and the results are illustrated in Figure 10. As the augment ratio initially increases from 0 to 1.0, we observe improvements across performance metrics. However, when the augment ratio is increased beyond 1.0, these improvements become marginal or even negative. This indicates that excessive enhancement of the graph structure can increase learning difficulty. The increasing disparity between the enhanced and original graph topologies may also negatively impact performance. This study reveals that a reasonable augment ratio $\epsilon$ benefits the model by improving its sensitivity to bandwidth constraints. However, excessively high $\epsilon$ values provide only slight improvements or can even degrade performance.
Appendix H Detailed Information on Used Assets
@@
  • Virne is a Virtual Network Embedding (VNE) algorithm library, offering many heuristic and machine learning-based methods for VNE. The source code can be accessed at https://github.com/GeminiLight/virne. It is licensed under Apache 2.0. In this work, we derived baseline implementations from Virne and developed our method using this library, specifically version 0.5.0.
  • SNDlib is a library for telecommunication network design, which offers a collection of realistic network system topologies. This library is open-source and available at https://sndlib.put.poznan.pl, although the specific licensing terms are not clearly stated. In our real-world network validation (see Appendix G.3), we utilize network topologies such as GEANT and BRAIN, which are from SNDlib.
All datasets and codebases are publicly accessible. They focus on networking and communication. They are not directly related to human identities and include no offensive content and bias.

### v2

@@ G.3 Scalability Analysis
Large-scale Network Validation. Similar to previous work (Wang et al., 2023), we generate a random Waxman topology (Waxman, 1988) with 500 nodes and nearly 13,000 links, named WX500. This mimics the modern large-scale cloud cluster, which is more challenging due to the increased complexity of the topology. Additionally, we increase the VNE size distribution to a uniform range from 2 to 20 nodes, and set the arrival rate of VN requests $\eta$ to 0.5. All other simulation and hyperparameter settings remain consistent with those outlined in Section 4.1. To adapt the pretrained CONAL model for WX100 to this larger scenario, we fine-tune it on WX500 via transfer learning. The pretraining on WX100 takes approximately 7.274 hours, and the fine-tuning stage on WX500 consumes an additional 4.326 hours. By leveraging transfer learning, we accelerate the training efficiency of CONAL, facilitating the rapid acquisition of a CONAL model suitable for large-scale scenarios. The results are illustrated in Figure 8. We observe that CONAL consistently outperforms the baselines in terms of both VN_ACR and LT_REV. CONAL also demonstrates superior performance in the LT_R2C, which is only marginally lower than NEA-VNE. Regarding the AVG_ST, NEA-VNE, MCTS-VNE, and PSO-VNE exhibit significantly higher time consumption, whereas CONAL maintains a competitive running efficiency similar to PG-CNN, A3C-GCN, and DDPG-ATT. Despite the increased complexity of the WX500 topology, CONAL maintains a balance between performance and computational efficiency. This analysis shows the scalability and efficiency of CONAL in large-scale network scenarios.
@@
Solving Time Scale Analysis. We investigate solving time across different physical network sizes to evaluate CONAL’s time scalability. Specifically, we increase the size of physical networks from 200 to 1,000 nodes with a step of 200 to simulate networks of varying scales. Due to enough link connectivity of large-scale network systems, we set the maximum path length $k$ to 4. Figure 9 shows that as the size increases, NRM-VNE, GRC-VNE and GAL-VNE show the most sightly increased average solving time. In contrast, PSO-VNE and GA-VNE experience rapid growth in solving time due to their reliance on extensive search. CONAL exhibits a similar trend with A3C-GCN and DDPG-ATT, since they are all based on constructive solution paradigms. As the size of the physical network increases, the solution time of these algorithms does not explode, still competitive with MCTS and NEA-VNE. Overall, the solving time of CONAL remains efficient even at larger network scales while offering great performance, making it a viable solution for real-time decision-making in large-scale environments.
G.4 Real-world Network Topology Validation
@@
G.5 Extension to Latency-aware VNE
Table 4: Experimental results on the variation of latency-aware VNE.
VN_RAC $\uparrow$ LT_R2C $\uparrow$ LT_REV (×$\times 10^{7}$) $\uparrow$
NRM-VNE (Fan et al., 2023) 0.655 $\pm$ 0.017 0.469 $\pm$ 0.005 7.168 $\pm$ 0.091
GRC-VNE (Gong et al., 2014) 0.454 $\pm$ 0.028 0.509 $\pm$ 0.008 5.281 $\pm$ 0.281
NEA-VNE (Fan et al., 2023) 0.714 $\pm$ 0.014 0.574 $\pm$ 0.006 8.186 $\pm$ 0.120
GA-VNE (Zhang et al., 2019) 0.717 $\pm$ 0.022 0.528 $\pm$ 0.004 7.712 $\pm$ 0.073
PSO-VNE (Jiang & Zhang, 2021) 0.709 $\pm$ 0.020 0.465 $\pm$ 0.003 7.457 $\pm$ 0.141
MCTS-VNE (Haeri & Trajković, 2017) 0.676 $\pm$ 0.017 0.489 $\pm$ 0.005 6.882 $\pm$ 0.102
PG-CNN (Zhang et al., 2022) 0.659 $\pm$ 0.019 0.500 $\pm$ 0.004 7.026 $\pm$ 0.146
DDPG-ATT (He et al., 2023a) 0.672 $\pm$ 0.024 0.477 $\pm$ 0.005 7.541 $\pm$ 0.096
A3C-GCN (Zhang et al., 2024b) 0.724 $\pm$ 0.020 0.552 $\pm$ 0.007 8.222 $\pm$ 0.125
GAL-VNE (Geng et al., 2023) 0.692 $\pm$ 0.028 0.567 $\pm$ 0.013 7.403 $\pm$ 0.345
CONAL 0.781 $\pm$ 0.027 0.604 $\pm$ 0.009 9.458 $\pm$ 0.161
In edge computing, latency is an additional quality of service (QoS) requirement, leading to a variant of VNE, known as latency-aware VNE. This variant incorporates latency constraints in addition to traditional VNE requirements. Specifically, a constraint is introduced for link mapping, where the total latency of the physical links in a path $p_{p}=f_{L}(l_{v})$ routing virtual link $l_{v}$ must not exceed the latency requirement of $l_{v}$. If any virtual link fails to meet its latency requirement, the VN request is rejected. Similar to previous studies (Jin et al., 2020; Adoga & Pezaros, 2023), we randomly assign coordinates to each physical node in the WX100 topology used in the main experiments, where each node represents a network edge. The latency for each physical link is then calculated based on the distance between paired physical nodes, and is normalized to a range of [0, 100] ms. For each VN request, the latency requirements of its virtual links are uniformly distributed within the range of [100, 500] ms. All other experimental settings remain consistent with those in the main experiments. To optimize these multi-type constraints, CONAL employs separate Lagrange multipliers for both resource and latency constraints, which are simultaneously managed by the neural lambda network $\Lambda$.
Table 4 presents the experimental results in latency-aware VNE. We observe that CONAL significantly outperforms all baseline methods, achieving the highest VN_RAC (0.781), LT_R2C (0.604), and LT_REV (9.458). Notably, the performance of GAL-VNE drops dramatically compared to the main experiment, even performing worse than the heuristic NEA-VNE. This is primarily because GAL-VNE relies on a one-shot solving paradigm and cannot effectively manage procedural delay constraints. In contrast, CONAL consistently maintains superior performance, thanks to its constraint-aware graph representation and adaptive optimization techniques, which balance both resource and latency constraints. These results highlight CONAL’s robust handling of complex edge computing scenarios for latency-aware VNE, making it an extensible solution for variations in diverse networking paradigms.
G.6 Hyperparameter Impact Study
@@
The impact of update interval of surrogate policy. This parameter $\mu$ controls the update interval of surrogate policy $\pi^{\prime}$ during training. We vary $\mu$ within the range [1, 5, 10, 20, 30] to explore its impact on key performance metrics, as shown in Figure 10. We observe that increasing $\mu$ initially leads to an improvement in all metrics. This suggests that extremely frequent updates of the surrogate policy make the budget values change rapidly, potentially leading to instability and divergence in the learning process. As $\mu$ increases beyond nearly 10, the improvements across these metrics tend to be stable and show minimal variation. This implies that while a moderate update interval enhances the model’s performance, too-slow updates do not offer a significant further improvement on performance and may even increase computational overhead for training.
The impact of the augment ratio. This parameter determines the proportion of links to be added based on the number of nodes in the network. We change the augment ratio $\epsilon$ within the range [0, 0.5, 1.0, 1.5, 2.0] and the results are illustrated in Figure 11. As the augment ratio initially increases from 0 to 1.0, we observe improvements across performance metrics. However, when the augment ratio is increased beyond 1.0, these improvements become marginal or even negative. This indicates that excessive enhancement of the graph structure can increase learning difficulty. The increasing disparity between the enhanced and original graph topologies may also negatively impact performance. This study reveals that a reasonable augment ratio $\epsilon$ benefits the model by improving its sensitivity to bandwidth constraints. However, excessively high $\epsilon$ values provide only slight improvements or can even degrade performance. Generally, setting $\epsilon$ = 1.0 or a value close to it provides a balanced trade-off between performance enhancement and model robustness.
Appendix H Used Benchmark and Assets
@@
  • Virne is a widely-adopted VNE benchmarking library, offering many heuristic and machine learning-based methods for VNE. The source code can be accessed at https://github.com/GeminiLight/virne. It is licensed under Apache 2.0. In this work, we derived baseline implementations from Virne and developed our method using this library.
  • SNDlib is a well-known library for telecommunication network design, which offers a collection of realistic network system topologies. This library is open-source and available at https://sndlib.put.poznan.pl, although the specific licensing terms are not clearly stated. In our real-world network validation (see Appendix G.4), we utilize network topologies such as GEANT and BRAIN, which are from SNDlib.